OpenStudy (anonymous):

solve: 2sin(5x+4)=-sqroot3

OpenStudy (anonymous):

sin(5x+4)=-sqrt(3) /2 -----> 5x+4 = 7pi/6 ---> x= ((7pi/6)-4)/5 or 5x+4=11pi/6 ----> x=((11pi/6)-4)/5

OpenStudy (anonymous):

t=5x+4 2sin(t)=-sqrt(3) sin(t)=-sqrt(3)/2 ==> t=2/3*pi + 2*pi*k 5x+4 = 2*pi/3 + 2*pi*k x=2*pi*k/5+2*pi/15-4/5

OpenStudy (anonymous):

the sine of an angle that equals -sqrt(3)/2 is 4pi/3 .. Not 7pi/6.---->

OpenStudy (anonymous):

-----> 5x+4 = 4pi/3 ---> x= ((4pi/3)-4)/5